OpenStudy (narissa):

determine if the statement is true or false when evaluated 2x^2-4x+1>2x-x^2/4 when x=2

OpenStudy (narissa):

@satellite73

OpenStudy (narissa):

isn't is also true?

OpenStudy (narissa):

@jim_thompson5910

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

guess we have to compute

OpenStudy (narissa):

hi

OpenStudy (fibonaccichick666):

so what did you try narissa?

OpenStudy (narissa):

the first part =9 right?

OpenStudy (misty1212):

put \(2\) in to \[2\times 2x^2-4x+1\] get \[2^2-4\times 2+1\]

OpenStudy (fibonaccichick666):

so first half is this yea? \[ 2x^2-4x+1\] In that case, how did you get 9?

OpenStudy (misty1212):

oops i made a typo

OpenStudy (misty1212):

\[2\times 2^2-4\times 2+1\]

OpenStudy (narissa):

2*2=4^=16-4*2=8+1=9

OpenStudy (misty1212):

oops not the correct order of operations

OpenStudy (fibonaccichick666):

ah, \(2^2=4\) right? so what when you compute you need to remember order of ops.

OpenStudy (misty1212):

square first, multiply second

OpenStudy (fibonaccichick666):

\(2(2^2)\) is a nice way to think about it

OpenStudy (narissa):

im confused then the statement is false

OpenStudy (fibonaccichick666):

well, let's figure it out first, so what do we get for the first part?

OpenStudy (narissa):

2*4=8

OpenStudy (fibonaccichick666):

ok, so what does the first part = when we do order of ops correctly?

OpenStudy (narissa):

confused

OpenStudy (fibonaccichick666):

ok, where at?

OpenStudy (narissa):

8-8+1 so it =1

OpenStudy (fibonaccichick666):

yes, that is true

OpenStudy (fibonaccichick666):

so now we need to do the second part

OpenStudy (narissa):

yaaa!

OpenStudy (narissa):

4-4=0

OpenStudy (narissa):

so its true

OpenStudy (fibonaccichick666):

so is this the second half? \[2x-\frac{x^2}{4}\]

OpenStudy (narissa):

2*2=4/4-4=0

OpenStudy (fibonaccichick666):

wait wait wait

OpenStudy (narissa):

ooops i think 3

OpenStudy (fibonaccichick666):

first, I need you to confirm the second half :)

OpenStudy (fibonaccichick666):

Is the eq I typed above accurate?

OpenStudy (narissa):

yes

OpenStudy (fibonaccichick666):

ok and yea, 3 seems a little better there, just to be clear, how did you get 3?

OpenStudy (narissa):

2*2=4 then the fraction =4/4 so 4-1 =3

OpenStudy (fibonaccichick666):

yea, now note the fraction is the same as division, so we are still following the order of ops here.

OpenStudy (fibonaccichick666):

now last step is 1>3 true?